# On the Adversarial Robustness of LASSO Based Feature Selection

### v1

@@ Abstract
In this paper, we present a novel feature selection method based on the Wasserstein distance. Feature selection plays a critical role in reducing the dimensionality of input data, thereby improving machine learning efficiency and generalization performance. Unlike traditional feature selection approaches that rely on criteria such as correlation or KL divergence, our method leverages the Wasserstein distance to measure the similarity between distributions of selected features and original features. This approach inherently accounts for similarities between classes, making it robust in scenarios involving noisy labels. Experimental results demonstrate that our method outperforms traditional approaches, particularly in challenging settings involving noisy labeled data.
@@
Feature selection, Wasserstein distance, label correlation, noisy label, graphical model.
 [^0] This work was accomplished at UC, Davis during Fuwei Li’s (e-mail: fuweily@gmail.com) Ph.D. study.
@@ I Introduction
Feature selection is a thoroughly studied topic in recent decades. In the preliminary phase of this research topic, it is aimed to increase the performance of the machine learning algorithm by selecting the most important features. At that time, machine learning algorithms are far less powerful than today’s, so by reducing the dimension of the input data features, the algorithm can learn much faster and more accurate. Also, the limited storage capacity at that time required the data samples to be as small as possible while keeping the most informative features. Now, we have much more powerful and efficient machine learning algorithms. For example, the deep neuron networks can handle billions of data samples. It seems that there is no need to do feature selection anymore. However, the data generated by humans and the Internet of Things(IoTs) have also increased dramatically. The rate grows even faster than the storage capacity as well as the computing power. In the near future when the IoTs play a more important part in our life, the data generated by IoTs and humans will be expected to grow exponentially.
@@
Algorithms of feature selection can be divided into three categories[3]: wrapper, embed, and filter methods. Wrapper methods use some exhaustive searching among different subsets of features and evaluate the feature sets based on a specific learning machine. These methods lead to very high computational complexity and are feasible only when the number of the features is small, it is seldom used nowadays.
The embed methods integrate feature selection into the design of the learning machine. Therefore, feature selection is accomplished since the training of the learning machine. The embed method includes various kinds of spars driven algorithms[4][5]. Since feature selection is done together with the training, they are time-saving. However, these methods require careful designing. Different learning algorithms and different parameters may lead to totally different features.
@@
Various kinds of criteria are proposed to select the best feature set, including the correlation based criterion[6], probability measure based criterion[7], and conditional likelihood based criterion[8]. Since most of the machine learning algorithms are intend to use a parametric model to mimic the empirical data distribution under some probabilistic distance measure, it is natural to choose the probability measure as the feature selection criterion. For a learning problem, given an instance of the feature $\mathbf{X}=\mathbf{x}$, we want to know the distribution of the classes. This class conditional distribution can further used to construct an optimal Bayesian estimator. So, the class conditional distribution is the most important part in a learning problem. After feature selection, we are only given a subset of the feature, $\mathbf{X}_{\theta}=\mathbf{x}_{\theta}$, we still intend to make the class distribution be similar to the original class distribution given the whole features. So, we minimize the distance between the two conditional distributions, the class conditional distribution given all the features $p(Y|\mathbf{X}=\mathbf{x})$ and the class conditional distribution given the selected features $p(Y|\mathbf{X}_{\theta}=\mathbf{x}_{\theta})$. Formally, we attempt to solve this minimization problem
$\underset{|\theta|=K}{\text{argmin}}\,D\,[\,p(Y|\mathbf{x}),\,p(Y|\mathbf{x}_{\theta})\,]$
where $D\,[\cdot,\cdot]$ is some probabilistic distance measure to tell the distance between two probability distributions, $|\theta|$ means the cardinality of the set $\theta$ and $K$ is the number of features we want to choose. Given the weights of different instances of feature $\mathbf{X}=\mathbf{x}$, it is better to use the expected value of $D$ as the distance measure. So, our minimization problem is
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}D\,[\,p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})\,]$ (1)
Later we will detail the choosing of our distance measure $D$ as the Wasserstein distance and explain why we choose that distance measure.
Traditionally, people may use measures induced by information theory, for example, the Kullback-Leibler (KL) divergence, f-divergence, and $\chi^{2}$ divergence. Among the three, the KL divergence is widely used[7][8][9]. Even though KL divergence is simple, it still has some weaknesses. The definition of KL divergence between $p(Y|\mathbf{x})$ and $p(Y|\mathbf{x}_{\theta})$ is $\sum_{i=1}^{n_{c}}p(y=i|\mathbf{x})\ln\frac{p(y=i|\mathbf{x})}{p(y=i|\mathbf{x}_{\theta})}$. It can also be written as an expectation form, that is, $\mathbb{E}_{p(Y|\mathbf{x})}\ln\frac{p(y|\mathbf{x})}{p(y|\mathbf{x}_{\theta})}$. The definition indicates the KL divergence treats all the classes equally. For example, if $p(Y|\mathbf{x})=[0.5,0.5]$ and $p(Y|\mathbf{x}_{\theta_{1}})=[0.1,0.9]$, and another distribution is $p(Y|\mathbf{x}_{\theta_{2}})=[0.9,0.1]$, then we have $D[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{1}})]=D[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{2}})]$. Even if we have two totally different distributions $p(Y|\mathbf{x}_{\theta_{1}})$ and $p(Y|\mathbf{x}_{\theta_{2}})$, the KL distance measure can not distinguish them.
@@
One reason why KL distance is so widely used is its simple and decomposable computation. With this decomposable computing advantage, we can design efficient machine learning algorithms. However, it fails to capture the relationship between different classes. Relationships between classes are very common in real practice. For example, in the text classification problem, one article is assigned to several classes which have hierarchical structures[10]. In ImageNet[11], some of the image labels have very high correlation. Several papers have used this correlation in classification problems[12][13][14]. Nevertheless, little work has done in the context of feature selection. In this work, we assume there is a ground metric, a prior, describing the distances between classes. Here is a simple example to show how useful of this inter-class information. Suppose there are three classes: cat, husky and alaska. If the distribution based on the full features is $p(Y|\mathbf{x})=[0,1,0]$ and two distributions based on different sub features $p(Y|\mathbf{x}_{\theta_{1}})=[1,0,0]$ and $p(Y|\mathbf{x}_{\theta_{2}})=[0,0,1]$. In this case, we would better choose $\mathbf{x}_{\theta_{2}}$ as the selected feature due to the fact that misclassifying a dog as a cat is more significant than as another kind of dog.
@@
Wasserstein distance is also called Earth Movement distance[15]. It is a probability distance measure which inherently uses the class distance. Wasserstein distance is used in many applications including computation graphics[16], clustering[17], Generative Adversarial Networks[18], and so forth. In the latter section, we will detail the computation of Wasserstein distance and how it is used in our feature selection problem.
@@
Another common problem many machine learning tasks encounter is the label noise. It is known that the label noise will dramatically decrease the classifier performance[19]. Sources of label noise may come from various ways. First, label noise may simply come from the transmission process. For example, if the class labeling work is distributed to the workers by the wireless communication channel, the communication noise may add to the labels[19]. Second, label noise may come from the lack of information. For instance, to tell whether a person is male or female, providing his/her voice data is not sufficient. In medical diagnosis, a doctor usually uses several reports provided by different inspection methods to determine a person has a certain kind of illness. Third, labels may come from some unreliable sources. In particular, now we have billions of data need to be labeled and the price to do label by human experts is very high. Instead, people are trying to use some cheap but not reliable source, like Amazon Mechanical Turk[20][21]. Fourth, when the class is subjective, for example, in medical diagnosis, for a certain medical image, different experts may give different results. This happens especially when two classes have some conceptual confusions. So, label noise commonly exists in real practical dataset.
@@
Various methods have been introduced to tackle the label noise problem[22]. For example, the filter based methods firstly detect the wrongly labeled data, correct them, and then use a traditional algorithm to do classification. Others use a specific label noise model to capture how label noise is introduced[23] [24] [25]. Even though so many label noise models are proposed, different models can only capture one certain aspect of the label noise. There is no unified way to model the label noise. In real practice, it is even more complicated, because in most cases, there is no ground truth label noise model. Though there is no unified way to describe the label noise, all of them admit that label noise is more likely to happen between two similar objects. We will show that our Wasserstein distance based algorithm captures this property naturally. Thus our algorithm is more stable and more robust to the label noise.
In the following section, we will first introduce the Wasserstein distance. Next, we will introduce the probabilistic graphic model and show how Markov Blanket has been used in the feature selection problem. Then, we will introduce our feature selection algorithm. Last, we will show the performance of our algorithm in the noisy label settings.
II The Wasserstein Distance
II-A Definition of Wasserstein Distance
@@
$D_{wass}(\nu_{1},\nu_{2})=\inf_{\pi\in\Pi(\nu_{1},\nu_{2})}\int_{\mathcal{C}_{1}\times\mathcal{C}_{2}}d(c_{1},c_{2})\pi(\mathrm{d}c_{1},\mathrm{d}c_{2})$ (2)
where $\Pi(\nu_{1},\nu_{2})$ is the set of joint probability measure with marginal probability measure $\nu_{1}$ and $\nu_{2}$ respectively. Solving this optimization problem means to find the joint probability which minimizes the cost of transporting one mass with measure $\nu_{1}$ to the corresponding mass with measure $\nu_{2}$. In this paper, we only focus on the discrete measure on the classes. In this case, the Wasserstein distance between probability measure $\mathbf{p}(Y)=[p(y_{1}),p(y_{2}),\cdots,p(y_{n_{c}})]$ and $\mathbf{q}(Y)=[q(y_{1}),q(y_{2}),\cdots,q(y_{c})]$ can be computed by solving the optimization problem:
@@
$Q_{i,j}\geq 0,\,\forall\,i,j\in\{1,2,\cdots,n_{c}\}$ (3)
where $\mathbf{Q}$ is the joint probability matrix with marginal distribution $\mathbf{p}(Y)=[p(Y=c_{1}),p(Y=c_{2}),\cdots,p(Y=c_{n})]$ and $\mathbf{q}(Y)=[q(Y=c_{1}),q(Y=c_{2}),\cdots,q(Y=c_{n})]$; $\mathbf{D}$ is the distance matrix with its elements defined as $D_{ij}=d(c_{i},c_{j})$ and $d(c_{i},c_{j})$ defines the distance from class $c_{i}$ to class $c_{j}$. The distance $d(c_{i},c_{j})$ measure the dissimilarity between class $c_{i}$ and class $c_{j}$. This is a linear programming problem. To solve this linear programming, the computational complexity would be $\mathcal{O}(n_{c}^{3.5})$[27]. The computation is heavy when the number of classes is large. Recently, [28] proposed a regularized form of this problem. They added a negative entropy of the joint distribution matrix regularization term in the objective function. Accordingly, we are solving the minimization problem with an objective function as
$\min:\text{tr}(\mathbf{Q}^{\top}\mathbf{D})-\frac{1}{\lambda}h(\mathbf{Q})$
@@
III Markov Blanket and Feature Selection
Markov Blanket is widely used in feature selection[29][30][31][32]. Markov Blanket is the smallest feature set which contains all the information about the target. In a directed probabilistic graphic model, the Markov Blanket of a node is the nodes’ parents, children and spouses. Figure(2) shows a probabilistic graphic model, where the Markov Blanket of node A contains the nodes in the shaded area. More formally, the Markov Blanket of the target $Y$ is the smallest set $\mathcal{B}\subset\{1,2,\cdots,M\}$ which have $p(Y=y|\mathbf{X}=\mathbf{x})=p(Y=y|\mathbf{X}_{\mathcal{B}}=\mathbf{x}_{\mathcal{B}})$ for every instance of $\mathbf{X}$, where the elements of $\mathcal{B}$ indicate which features are chosen.

### v3

# Feature Selection Based on Wasserstein Distance

@@ Abstract
This paper presents a novel feature selection method leveraging the Wasserstein distance to improve feature selection in machine learning. Unlike traditional methods based on correlation or Kullback-Leibler (KL) divergence, our approach uses the Wasserstein distance to assess feature similarity, inherently capturing class relationships and making it robust to noisy labels. We introduce a Markov blanket-based feature selection algorithm and demonstrate its effectiveness. Our analysis shows that the Wasserstein distance-based feature selection method effectively reduces the impact of noisy labels without relying on specific noise models. We provide a lower bound on its effectiveness, which remains meaningful even in the presence of noise. Experimental results across multiple datasets demonstrate that our approach consistently outperforms traditional methods, particularly in noisy settings.
@@
Feature selection, Wasserstein distance, label correlation, noisy label, probabilistic graphical model.
 [^0] This work was accomplished at UC, Davis, during Fuwei Li’s (e-mail: fuweily@gmail.com) Ph.D. study under the supervision of Lifeng Lai and Shuguang Cui.
@@ I Introduction
Feature selection is a thoroughly studied topic in recent decades. In the preliminary phase of this research topic, it is aimed to increase the performance of the machine learning algorithm by selecting the most important features. At that time, machine learning algorithms are far less powerful than today’s, so by reducing the dimension of the input data features, the algorithm can learn much faster and more accurate. Also, the limited storage capacity at that time required the data samples to be as small as possible while keeping the most informative features. Now, we have much more powerful and efficient machine learning algorithms. For example, the deep neuron networks can handle billions of data samples. It seems that there is no need to do feature selection anymore. However, the data generated by humans and Internet of things(IoTs) have also increased dramatically. The rate grows even faster than the storage capacity as well as the computing power. In the near future when the IoTs play a more important part in our life, the data generated by IoTs and humans will be expected to grow exponentially.
@@
Algorithms of feature selection can be divided into three categories[3]: wrapper, embed, and filter methods. Wrapper methods use some exhaustive searching among different subsets of features and evaluate the feature sets based on a specific learning machine. These methods lead to very high computational complexity and is feasible only when the number of the features is small, it is seldom used nowadays.
The embed methods integrate feature selection into the design of the learning machine. Therefore, feature selection is accomplished since the training of the learning machine. The embed method includes various kinds of sparse driven algorithms[4][5]. Since feature selection is done together with the training, they are time-saving. However, these methods require careful designing. Different learning algorithms and different parameters may lead to totally different features.
@@
Within filter methods, various kinds of criteria are proposed to select the best feature set, including the correlation based criterion[6], probability measure based criterion[7], and conditional likelihood based criterion[8]. Since most of the machine learning algorithms are intend to use a parametric model to mimic the empirical data distribution under some probabilistic distance measure, it is natural to choose the probability measure as the feature selection criterion.
Traditionally, people may use measures induced by information theory, for example, the Kullback-Leibler (KL) divergence, f-divergence, and $\chi^{2}$ divergence. Among the three, the KL divergence is widely used[7][8][9]. Even though KL divergence is simple, it still has some weaknesses. The definition of KL divergence between $p(Y|\mathbf{x})$ and $p(Y|\mathbf{x}_{\theta})$ is $\sum_{i=1}^{n_{c}}p(y=i|\mathbf{x})\ln\frac{p(y=i|\mathbf{x})}{p(y=i|\mathbf{x}_{\theta})}$, where $n_{c}$ is the number of classes, $Y$ is label of the class, $\mathbf{x}$ is the whole feature set before feature selection, and $\mathbf{x}_{\theta}$ is a subset of feature indexed by $\theta$. The definition indicates the KL divergence treats all the classes equally. For example, if $p(Y|\mathbf{x})=[0.5,0.5]$ and $p(Y|\mathbf{x}_{\theta_{1}})=[0.1,0.9]$, and another distribution is $p(Y|\mathbf{x}_{\theta_{2}})=[0.9,0.1]$, then we have $KL[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{1}})]=KL[p(Y|\mathbf{x}),p(Y|\mathbf{x}_{\theta_{2}})]$. Even if we have two totally different distributions $p(Y|\mathbf{x}_{\theta_{1}})$ and $p(Y|\mathbf{x}_{\theta_{2}})$, the KL distance measure can not distinguish them.
@@
One reason why KL distance is so widely used is its simple and decomposable computation. With this decomposable computing advantage, we can design efficient machine learning algorithms. However, it fails to capture the relationship between different classes. Relationships between classes are very common in real practice. For example, in the text classification problem, one article is assigned to several classes which have hierarchical structures[10]. In ImageNet[11], some of the image labels have very high correlation. Several papers have used this correlation in classification problems[12][13][14]. Nevertheless, little work has done in the context of feature selection. In this work, we assume there is a ground metric, a prior, describing the dissimilarities between classes. Fig.1 gives a simple example to show how useful of this inter-class information is. Suppose there are three classes: cat, husky and alaska. If the distribution based on the full features is $p(Y|\mathbf{x})=[0,1,0]$ and two distributions based on different sub features $p(Y|\mathbf{x}_{\theta_{1}})=[1,0,0]$ and $p(Y|\mathbf{x}_{\theta_{2}})=[0,0,1]$. In this case, we would better choose $\mathbf{x}_{\theta_{2}}$ as the selected feature due to the fact that confusing a dog with a cat is more severe than confusing with two kinds of dogs. However, if we ignore the class similarity and use KL divergence, it can not tell which subset of feature is better. While, Wasserstein distance is a natural probabilistic distance measure which inherently use the class similarities.
@@
Wasserstein distance is also called Earth Movement distance[15]. Wasserstein distance is used in many applications including computation graphics[16], clustering[17], Generative Adversarial Networks[18], and so forth. In the latter section, we will detail the computation of Wasserstein distance and how it is used in our feature selection problem.
@@
Another common problem many machine learning tasks encounter is the label noise. It is known that the label noise will dramatically decrease the classifiers’ performance[19]. Sources of label noise may come from various ways. First, label noise may simply come from the transmission process. For example, if the class labeling work is distributed to the workers by the wireless communication channel, the communication noise may add to the labels[19]. Second, label noise may come from the lack of information. For instance, to tell whether a person is male or female, providing his/her voice data is not sufficient. In medical diagnosis, a doctor usually uses several reports provided by different inspection methods to determine a person has a certain kind of illness. Third, labels may come from some unreliable sources. In particular, now we have billions of data need to be labeled and the price to do label by human experts is very high. Instead, people are trying to use some cheap but not reliable source, like Amazon Mechanical Turk[20][21]. Fourth, when the class is subjective, for example, in medical diagnosis, for a certain medical image, different experts may give different results. This happens especially when two classes have some conceptual confusions. So, label noise commonly exists in real practical dataset.
@@
Various methods have been introduced to tackle the label noise problem[22]. For example, the filter based methods firstly detect the wrongly labeled data, correct them, and then use a traditional algorithm to do classification. Others use a specific label noise model to capture how label noise is introduced[23] [24] [25]. Even though so many label noise models are proposed, different models can only capture one certain aspect of the label noise. There is no unified way to model the label noise. In real practice, it is even more complicated, because in most cases, there is no ground truth label noise model. Though there is no unified way to describe the label noise, all of them admit that label noise is more likely to happen between two similar objects. We will show that our Wasserstein distance based algorithm captures this property naturally. Thus our algorithm is more robust to the label noise.
The contribution of this work can summarized as follows:
  • We are the first to use Wasserstein distance in the context of feature selection. In the existing works, people tended to use information based criteria to choose the most important features. Among which, KL divergence is the most widely used criterion. However, it lacks the ability to capture the correlation between classes. Wasserstein distance has the inherent nature to capture the similarities between classes. Thus makes it more suitable to more commonly seen feature selection problem. To show the Wasserstein distance is suitable for feature selection problem, we prove several useful Wasserstein distance induce independence properties. With the facility of the Wasserstein distance, we proposed a Markov blanked based feature selection algorithm and analyzed the computational complexity of the algorithm.
  • We analyzed the performance under the noisy labeled settings and showed that Wasserstein distance based feature selection has the potential to reduce the impact of noise labels. Noisy labels exists in many practical feature selection problems and how the noisy labels were introduced is every complex. Directly model the noisy label is impractical. Our analysis does not depend on a specific noisy model. With the facility of the Wasserstein distance, we give a lower bound on how good it is using the Wasserstein distance to do feature selection even some noisy labels exist. Also, we show that in general this bound is small and meaningful under the Wasserstein distance.
  • We do several simulations, which includes text analysis, email classification, and image classification, to show our proposed Wasserstein distance based feature selection algorithm is useful and superior to its corresponding KL distance based feature selection algorithm.
The outline of the subsequent parts is: we formulate the feature selection problem rigorously, introduce the Wasserstein distance, give a brief introduction to Markov Blanket, show the independence properties of Wasserstein distance, and introduce our algorithm; analyze the performance of our algorithm in the noisy label settings under the Wasserstein distance; demonstrate the performances under serveral simulations.
II The Algorithm
II-A Problem Formulation
For a learning problem, given an instance of the feature $\mathbf{X}=\mathbf{x}$, we want to know the distribution of the classes. This class conditional distribution can further used to construct an optimal Bayesian estimator. So, the class conditional distribution is most important in a learning problem. After feature selection, we are only given a subset of the feature, $\mathbf{X}_{\theta}=\mathbf{x}_{\theta}$, we still intend to make the class distribution be similar to the original class distribution given the whole features. So, we minimize the distance between the two conditional distributions, the class conditional distribution given all the features $p(Y|\mathbf{X}=\mathbf{x})$ and the class conditional distribution given the selected features $p(Y|\mathbf{X}_{\theta}=\mathbf{x}_{\theta})$. Formally, we attempt to solve this minimization problem
$\underset{|\theta|=K}{\text{argmin}}\,M\,[\,p(Y|\mathbf{x}),\,p(Y|\mathbf{x}_{\theta})\,],$
where $M\,[\cdot,\cdot]$ is some probabilistic distance measure to tell the distance between two probability distributions, $|\theta|$ means the cardinality of the set $\theta$ and $K$ is the number of features we want to choose. Given the weights of different instances of feature $\mathbf{X}=\mathbf{x}$, it is better to use the expected value of $M$ as the distance measure. So, our minimization problem is
$\underset{|\theta|=K}{\text{argmin}}\,\mathbb{E}_{\mathbf{X}}M\,[\,p(Y|\mathbf{X}),\,p(Y|\mathbf{X}_{\theta})\,],$ (1)
In this paper, we choose $M$ as the Wasserstein distance. We will explain the details of the Wasserstein distance in the following.
II-B The Wasserstein Distance
@@
$W(\nu_{1},\nu_{2})=\inf_{\pi\in\Pi(\nu_{1},\nu_{2})}\int_{\mathcal{C}_{1}\times\mathcal{C}_{2}}d(c_{1},c_{2})\pi(\mathrm{d}c_{1},\mathrm{d}c_{2}),$ (2)
where $\Pi(\nu_{1},\nu_{2})$ is the set of joint probability measure with marginal probability measure $\nu_{1}$ and $\nu_{2}$ respectively. Solving this optimization problem means to find the joint probability which minimizes the cost of transporting one mass with measure $\nu_{1}$ to the corresponding mass with measure $\nu_{2}$. In this paper, we only focus on the discrete measure on the classes. In this case, the Wasserstein distance between probability measure $\mathbf{p}(Y)=[p(y_{1}),p(y_{2}),\cdots,p(y_{n_{c}})]$ and $\mathbf{q}(Y)=[q(y_{1}),q(y_{2}),\cdots,q(y_{c})]$ can be computed by solving the linear optimization problem:
@@
$Q_{i,j}\geq 0,\,\forall\,i,j\in\{1,2,\cdots,n_{c}\},$ (3)
where $\mathbf{Q}$ is the joint probability matrix with marginal distributions be $\mathbf{p}(Y)=[p(Y=c_{1}),p(Y=c_{2}),\cdots,p(Y=c_{n})]$ and $\mathbf{q}(Y)=[q(Y=c_{1}),q(Y=c_{2}),\cdots,q(Y=c_{n})]$; $\mathbf{D}$ is the distance matrix with its elements defined as $D_{ij}=d(c_{i},c_{j})$ and $d(c_{i},c_{j})$ defines the distance from class $c_{i}$ to class $c_{j}$ which measures the dissimilarity between class $c_{i}$ and class $c_{j}$. This is a linear programming problem. To solve this linear programming, the computational complexity would be $\mathcal{O}(n_{c}^{3.5})$[27]. The computation is heavy when the number of classes is large. Recently, [28] proposed a regularized form of this problem. They added a negative entropy of the joint distribution matrix regularization term in the objective function. Accordingly, we are solving the minimization problem with an objective function as
$\min:\text{tr}(\mathbf{Q}^{\top}\mathbf{D})-\frac{1}{\lambda}H(\mathbf{Q}),$
@@
II-C Markov Blanket
Markov Blanket is widely used in feature selection[29][30][31][32]. Markov Blanket is the smallest feature set which contains all the information about the target. In a directed probabilistic graphic model, the Markov Blanket of a node is the nodes’ parents, children and spouses. Fig.2 shows a probabilistic graphic model, where the Markov Blanket of node A contains the nodes in the shaded area. More formally, the Markov Blanket of the target $Y$ is the smallest set $\mathcal{B}\subset\{1,2,\cdots,M\}$ which have $p(Y=y|\mathbf{X}=\mathbf{x})=p(Y=y|\mathbf{X}_{\mathcal{B}}=\mathbf{x}_{\mathcal{B}})$ for every instance of $\mathbf{X}$.